Find the roots of each quadratic equation without making a table or a graph.
y=x^2+4x-5

Answer by scott8148(6628)

factoring ___ y = (x+5)(x-1)

0 = (x+5)(x-1)

0 = x + 5 ___ -5 = x

0 = x - 1 ___ 1 = x

roots are -5 and 1